# FunAudio-ASR Technical Report

### v2

@@ 6.2.2 Streaming ASR Performance
In order to evaluate the streaming ability of our FunAudio-ASR model, we evaluate the performance on the same test set used when evaluating the offline speech recognition ability of our FunAudio-ASR model. Table 3 lists the testing results. When comparing with Seed-ASR  (Bai et al., 2024), our FunAudio-ASR model have better performance at different test sets and test scenario.
Test set Seed-ASR FunAudio-ASR In-house 8.64 7.00 Fairfield 5.51 5.33 Home Scenario 9.7 5.33 Complex Background 15.48 12.50 English General 18.78 14.74 OpenSouce Test Sets 3.80 3.60
@@ 6.2.3 Evaluation on Noise Robustness
FunAudio-ASR Environment w/o NRT w/ NRT NRT + RL canteen 20.67 20.34 19.88 dinner 14.02 9.88 9.55 meeting 6.45 6.27 6.24 office 15.02 11.58 11.42 outdoor 10.12 9.85 9.58 park 13.67 11.88 11.37 shop 12.22 11.48 11.24 street 12.05 10.58 10.86 subway 14.11 13.31 13.29 supermarket 14.27 8.81 8.75 walkstreet 13.89 13.87 13.94 Average 13.32 11.58 11.45
@@ 6.2.5 Evaluation on Hotword Customization
For the hotword evaluation, we choose the audios with some special topics, including biology, math, religion, food, name, astronomy, chemistry, philosophy, and physics, as the recognition of the technical terms is crucial but still challenging for most ASR systems. Results in Table 6 shows that FunAudio-ASR can benefit from the hotword customization. On most topics, the recall rate can raise to more than 0.97 for FunAudio-ASR. The FunAudio-ASR shows good performance on the name topic, the recall can increase from 0.75 to 1.0. This indicates the hotword customization can really inspire the target keyword, rather than just provide contextual information.
@@ 6.2.6 Multilingual ASR Results
We also evaluate our multilingual ASR model FunAudio-ASR-ML on several open source test sets and in-house industry test sets. Table 7 lists the testing results. From Table 7, we can see that, on the Chinese and English open source test sets and in-house industry test sets, our multi-lingual ASR model FunAudio-ASR-ML have better or comparable effects when comparing with Kimi-Audio  (KimiTeam et al., 2025). We also compared our model with other multi-lingual ASR models, such as Whisper large v3 (Radford et al., 2023), dolphin-small  (Meng et al., 2025), and seamless-m4t large v2  (Communication et al., 2023). When comparing with these models, our FunAudio-ASR-ML model can also get the SOTA performance.
Language Test set Kimi-Audio Whisper Large v3 dolphin-small seamless-m4t-large-v2 FunAudio-ASR-ML fleurs 2.69 4.71 5.46 5.15 3.0 commonvoice 7.21 12.61 9.94 10.76 5.76 Chinese wenetspeech-test-net 5.37 9.83 9.63 9.87 6.48 aishell2-ios-test 2.56 4.83 4.37 4.79 2.60 in-house-test-set 36.42 16.54 9.67 14.85 7.91 fleus 4.4 4.11 N.A. 6.59 3.18 commonvoice 10.31 9.66 N.A. 7.63 7.67 English librispeech-test-clean 1.28 2.56 N.A. 2.56 1.62 librispeech-test-other 2.42 4.34 N.A. 4.84 3.39 in-house-test-set 12.40 11.78 N.A. 43.74 11.19 fleurs N.A. 6.07 15.86 9.36 8.09 Indonesian commonvoice N.A. 7.27 8.91 6.1 4.19 gigaspeech2-test N.A. 19.11 26.56 22.3 16.18 in-house-test-set N.A. 23.19 40.16 24.41 18.83 fleurs N.A. 8.48 9.66 9.25 7.04 Thai commonvoice N.A. 5.92 3.04 2.81 1.44 gigaspeech2-test N.A. 19.35 19.15 21.7 16.6 fleurs N.A. 6.51 15.62 8.07 6.33 Vietnamese commonvoice N.A. 13.51 12.73 13.85 13.49 gigaspeech2-test N.A. 13.82 31.98 43.31 8.66 in-house-test-set N.A. 11.46 40.82 32.10 6.9
@@ 6.2.7 Effect of Reinforcement Learning
Table 8 shows that RL plays a crucial role in FunAudio-ASR training, bringing approximately 4.1% and 9.2% relative improvement under offline and streaming conditions, respectively. For offline ASR, the performance gain is more pronounced on audio from noisy and complex environments compared to clean or open-source data. Notably, the improvement is even more significant in the streaming ASR setting. RL helps suppress both insertion and deletion errors, which may be attributed to early termination or premature prediction before full pronunciation.
@@
As shown in Table 6, RL effectively enhances hotword integration, leading to improvements in both accuracy and recall across most test sets. In certain domains, such as philosophy and religion, the RL model may achieve slightly lower accuracy or recall compared to the baseline; however, the overall WER still decreases. This is because, during RL training, keywords are selected based on the actual transcriptions rather than the input prompts, enabling FunAudio-ASR to better recognize domain-specific terminology—even for professional terms not explicitly included in the hotword list.
@@ 7 Limitations and Future Plans
Despite strong results across diverse evaluations, our FunAudio-ASR model still has some limitations. First, it is primarily optimized for Chinese and English—particularly for streaming performance and hotword customization—so support for other languages remains limited. Second, the effective context window is constrained; without an external voice activity detection (VAD) module, the system struggles to handle long-duration recordings robustly. Third, the current release does not support far-field or multi-channel audio. We plan to address these limitations in future work.
@@ 8 Conclusion
In this paper, we present FunAudio-ASR, a large-scale, LLM-based automatic speech recognition (ASR) system that leverages massive data, extensive model capacity, seamless integration of LLMs, and refinement learning to achieve state-of-the-art performance across diverse and challenging scenarios. Designed with practical deployment in mind, FunAudio-ASR incorporates key optimizations for real-world applications, including enhanced streaming capabilities, robustness to noise, effective handling of code-switching, and customizable hotword support. Experimental results reveal that while many LLM-based ASR systems perform well on open-source benchmarks, they often underperform on real-world industrial evaluation sets. In contrast, FunAudio-ASR—through production-oriented design and optimization—demonstrates superior accuracy on practical application datasets, establishing a new benchmark for high-performance, deployable ASR systems.

### v3

@@ 6.2.2 Streaming ASR Performance
In order to evaluate the streaming ability of our Fun-ASR model, we evaluate the performance on the same test set used when evaluating the offline speech recognition ability of our Fun-ASR model. Table 3 lists the testing results. When comparing with Seed-ASR  (Bai et al., 2024), our Fun-ASR model have better performance at different test sets and test scenario.
Test set Seed-ASR Fun-ASR Fun-ASR-nano In-house 8.64 7.00 7.97 Fairfield 5.51 5.33 6.92 Home Scenario 9.7 5.33 6.51 Complex Background 15.48 12.50 14.83 English General 18.78 14.74 16.70 OpenSouce Test Sets 3.80 3.60 5.13
@@ 6.2.3 Evaluation on Noise Robustness
Fun-ASR Environment w/o NRT w/ NRT NRT + RL canteen 20.67 20.34 19.88 dinner 14.02 9.88 9.55 meeting 6.45 6.27 6.24 office 15.02 11.58 11.42 outdoor 10.12 9.85 9.58 park 13.67 11.88 11.37 shop 12.22 11.48 11.24 street 12.05 10.58 10.86 subway 14.11 13.31 13.29 supermarket 14.27 8.81 8.75 walkstreet 13.89 13.87 13.94 Average 13.32 11.58 11.45
@@ 6.2.5 Evaluation on Hotword Customization
For the hotword evaluation, we choose the audios with some special topics, including biology, math, religion, food, name, astronomy, chemistry, philosophy, and physics, as the recognition of the technical terms is crucial but still challenging for most ASR systems. Results in Table 6 shows that Fun-ASR can benefit from the hotword customization. On most topics, the recall rate can raise to more than 0.97 for Fun-ASR. The Fun-ASR shows good performance on the name topic, the recall can increase from 0.75 to 1.0. This indicates the hotword customization can really inspire the target keyword, rather than just provide contextual information.
@@ 6.2.6 Multilingual ASR Results
We also evaluate our multilingual ASR model Fun-ASR-ML on several open source test sets and in-house industry test sets. Table 7 lists the testing results. From Table 7, we can see that, on the Chinese and English open source test sets and in-house industry test sets, our multi-lingual ASR model Fun-ASR-ML have better or comparable effects when comparing with Kimi-Audio  (KimiTeam et al., 2025). We also compared our model with other multi-lingual ASR models, such as Whisper large v3 (Radford et al., 2023), dolphin-small  (Meng et al., 2025), and seamless-m4t large v2  (Communication et al., 2023). When comparing with these models, our Fun-ASR-ML model can also get the SOTA performance.
Language Test set Kimi-Audio Whisper Large v3 dolphin-small seamless-m4t-large-v2 Fun-ASR-ML fleurs 2.69 4.71 5.46 5.15 3.0 commonvoice 7.21 12.61 9.94 10.76 5.76 Chinese wenetspeech-test-net 5.37 9.83 9.63 9.87 6.48 aishell2-ios-test 2.56 4.83 4.37 4.79 2.60 in-house-test-set 36.42 16.54 9.67 14.85 7.91 fleus 4.4 4.11 N.A. 6.59 3.18 commonvoice 10.31 9.66 N.A. 7.63 7.67 English librispeech-test-clean 1.28 2.56 N.A. 2.56 1.62 librispeech-test-other 2.42 4.34 N.A. 4.84 3.39 in-house-test-set 12.40 11.78 N.A. 43.74 11.19 fleurs N.A. 6.07 15.86 9.36 8.10 Indonesian commonvoice N.A. 7.27 8.91 6.1 5.49 gigaspeech2-test N.A. 19.11 26.56 22.3 16.71 in-house-test-set N.A. 23.19 40.16 24.41 18.36 fleurs N.A. 8.48 9.66 9.25 6.0 Thai commonvoice N.A. 5.92 3.04 2.81 0.9 gigaspeech2-test N.A. 19.35 19.15 21.7 19.09 fleurs N.A. 6.51 15.62 8.07 5.50 Vietnamese commonvoice N.A. 13.51 12.73 13.85 7.41 gigaspeech2-test N.A. 13.82 31.98 43.31 8.98 in-house-test-set N.A. 11.46 40.82 32.10 7.06
@@ 6.2.7 Effect of Reinforcement Learning
Table 8 shows that RL plays a crucial role in Fun-ASR training, bringing approximately 4.1% and 9.2% relative improvement under offline and streaming conditions, respectively. For offline ASR, the performance gain is more pronounced on audio from noisy and complex environments compared to clean or open-source data. Notably, the improvement is even more significant in the streaming ASR setting. RL helps suppress both insertion and deletion errors, which may be attributed to early termination or premature prediction before full pronunciation.
@@
As shown in Table 6, RL effectively enhances hotword integration, leading to improvements in both accuracy and recall across most test sets. In certain domains, such as philosophy and religion, the RL model may achieve slightly lower accuracy or recall compared to the baseline; however, the overall WER still decreases. This is because, during RL training, keywords are selected based on the actual transcriptions rather than the input prompts, enabling Fun-ASR to better recognize domain-specific terminology—even for professional terms not explicitly included in the hotword list.
@@ 7 Limitations and Future Plans
Despite strong results across diverse evaluations, our Fun-ASR model still has some limitations. First, it is primarily optimized for Chinese and English—particularly for streaming performance and hotword customization—so support for other languages remains limited. Second, the effective context window is constrained; without an external voice activity detection (VAD) module, the system struggles to handle long-duration recordings robustly. Third, the current release does not support far-field or multi-channel audio. We plan to address these limitations in future work.
@@ 8 Conclusion
In this paper, we present Fun-ASR, a large-scale, LLM-based automatic speech recognition (ASR) system that leverages massive data, extensive model capacity, seamless integration of LLMs, and refinement learning to achieve state-of-the-art performance across diverse and challenging scenarios. Designed with practical deployment in mind, Fun-ASR incorporates key optimizations for real-world applications, including enhanced streaming capabilities, robustness to noise, effective handling of code-switching, and customizable hotword support. Experimental results reveal that while many LLM-based ASR systems perform well on open-source benchmarks, they often underperform on real-world industrial evaluation sets. In contrast, Fun-ASR—through production-oriented design and optimization—demonstrates superior accuracy on practical application datasets, establishing a new benchmark for high-performance, deployable ASR systems.